# chattersPwCHparticipants with chatbot history

### v1

@@ 2.2. Setting and Participants
Prior work found ChatGPT to struggle with identifying issues in code, especially for logic errors (Tian et al., 2024; Hellas et al., 2023). For the purposes of generating meaningful student-chatbot interactions, it was important that the task could not be trivially solved by ChatGPT, but was at the same time attainable for the students. We thus designed an exercise that would be particularly challenging for the , which we confirmed in initial testing, without increasing the complexity or context size beyond what the students were accustomed to.
The exercise is presented in Figure 2. It consisted of two problems. In P1, participants should identify and trigger a null-related bug in a provided class. In P2, participants were then asked to add the missing null check to the original code, to prevent such errors. Participants were not allowed to reference other course material, look up information online, or ask other students for help during the experiment.
The experiment was conducted in-person in a controlled environment over the course of 13 sessions between November and December 2023. All participants were students of the same course. To accommodate the students’ schedules, they could participate immediately following each of their respective exercise groups during the week. Students of the course were informed of the study during a course lecture as well as through a Moodle announcement. Additionally, students were approached for recruitment in-person during their exercise groups. As incentive, all participants who completed the experiment also entered a gift card raffle.
Participation was completely voluntary and did not have any impact on the actual course grading. Participants were told the experiment was expected to take around 20 minutes, though no hard time limit was given. While participants were not misled about the purpose of the study, it was kept intentionally vague to reduce potential subject expectancy effects. It was described to them as a study of tool support for learning and problem-solving.
@@ 3.1. Task-Solving Evaluation (RQ1)
Figure 4. Performance in writing (question-answering) exercises.
@@ 3.1.1. Coding Exercises
Both GPT models produced syntactically correct code in all responses. Responses from GPT-3.5 had slightly more logical errors than GPT-4, and failed to meet the exercise requirements more than twice as often. Inconsistencies between the described intent and the actual implementation were rare, though in one particularly obvious case, GPT-3.5 provided multiple line comments describing logic that was entirely absent from the code.
Despite not being explicitly instructed to do so, large majority of the responses were aligned with what a student of the course would already know at that time.
Adherence to the course’s coding conventions was low overall. Neither language model’s baseline code generation – i.e. with no existing code provided as a reference – was in alignment with the course’s formatting guidelines. Even with context code, the majority of generated code did not follow the demonstrated bracketing style, nor were documentation comments added consistently.
@@ 3.2. Chatbot Use (RQ2)
3.2.1. Experiment Data
In total, 42 students participated in the experiment. 3 students participated in an initial test run with an older version of the exercise; their results were not included in the analysis. Additionally, two students ended the experiment after a short period (4.5 min. and 9 min.) without completing the tasks. We thus concluded that they dropped the participation and discarded their data. This leaves 37 participants for the analysis.
62.2 % of participants completed the entire exercise correctly, and 78.4 % successfully solved at least one problem. The average duration of experiment sessions was 941.9s (M = 859s, $\sigma$ = 559.2s) for students who completed both problems, and 1916.9s (M = 1784s, $\sigma$ = 493.7s) for those who did not. Across the experiment, 556 submission attempts were recorded, the vast majority of which (89 %) were for problem P1. Each participant made an average of 13.38 submission attempts for P1 (M = 12, $\sigma$ = 9.5), of which 67.5% resulted in compilation errors and 22.4% in a test failure because their code ran without errors. For P2, participants made an average of 1.65 attempts (M = 1, $\sigma$ = 2.32), of which 27.9% resulted in compilation errors and 13.1% failed because their code did not handle the error case.
Before starting the exercise, we asked participants to rate how challenging they found the course on a 5-point scale, with 1 being very easy and 5 being very difficult. We observed a moderate negative correlation between perceived course difficulty and exercise completion ($r=-0.47,p=0.003$), as well as a moderate positive correlation between perceived difficulty and experiment duration ($r=0.4,p=0.013$). This indicates that the students’ performance on the exercise aligned with their general experience in the course.
3.2.2. Usage of GAI
Before the exercise, participants were asked how often they use tools such as ChatGPT or GitHub Copilot. 16.2 % reported to have never used such tools before. 48.6 % reported using tools at least monthly (monthly through weekly or more than once a week). We observed a moderate positive correlation between the reported use of tools and the frequency of chatbot interactions during the experiment ($r=0.35,p=0.03$). We did not observe any significant relationships between the students’ reported use and their performance on the exercise.
@@ 3.2.2. Usage of GAI
The average experiment duration varied significantly between students who did and did not use the chatbot ($t=3.3,p=0.003$). For , the average duration was 26:15 (M = 26:03, $\sigma$ = 10:38). For users who did not use the chatbot, this was 14:36 (M = 10:30, $\sigma$ = 10:27). To account for participants leaving at arbitrary times if they could not solve the exercise within the first 20 minutes, we additionally separated them by exercise completion. When comparing the duration only among completed attempts, the difference is not significant ($t=1.9,p>0.05$). We therefore do not find a relationship between chatbot use and the time to arrive at the correct solution.
@@ 3.2.3. Student Prompts
In this section, all values are relative to the participants who used the chatbot during the experiment. Figure 5 shows the total number of prompts for each type, as well as the number of participants that have submitted such a prompt at least once. Overall, we found that students submitted more Codegen prompts than Support,with a ratio of 62% to 38% [^2] Others prompts are excluded from the totals here, as they were exclusively either off-topic or erroneously submitted..
We observed Solve prompts from most of the participants, as 73.9 % submitted a Solve prompt at some point during the experiment. Interestingly, we also found that some participants submitted more than two Solve prompts. Overall, the most common prompt type was Fix, though this was driven in part by two outliers, who each submitted 13 Fix prompts during their session. The least common prompt type was Explain, observed in only 39.1 % of conversations, constituting 6.8 % of prompts. Figure 6 shows the number of occurrences of each prompt type per-participant.
Figure 5. Total number of prompts by type (top) and number of participants that submitted at least one prompt of that type (bottom). Opacity indicates self-reported usage, with the highest opacity being “more than once a week”.
@@
Table 4. Correlation analysis of prompt labels to preceding survey responses using Kendall’s $\tau$. For individual labels, significance was adjusted by Bonferroni correction ($\alpha=0.05/5=0.01$).
@@ 3.2.4. Bot Responses
61 % of the 172 chatbot responses contained an auto-generated solution to one of the exercise problems, of which 97.1 % also contained a written explanation of the solution attempt. In total, 83.1 % of responses included an explanation, either for a generated solution, for an example code snippet, or for a general concept.
We found that the GPT-3.5 model was unable to solve problem P1 reliably. Out of 80 generated solutions, only 8.8 % were correct. The model performed better on P2, where 76.9 % of 13 generated solutions were correct.
88.1 % of the Inform responses were correct, indicating that when requesting only general information about basic programming concepts, GPT-3.5 does not appear to “hallucinate”. These included questions about when null-related errors occur in Java and how to produce one. Notably, one participant requested the documentation for the class NullPointerException, and the generated response correctly relayed it verbatim.
@@ 3.3.1. Similarity of Consecutive Submissions
Among submissions where participants had at least one Codegen response available to them, the average similarity to the generated code was 0.65 on P1 (M = 0.68, $\sigma$ = 0.26) and 0.66 on P2 (M = 0.71, $\sigma$ = 0.29). These numbers align with findings by Kazemitabaar et al. (Kazemitabaar et al., 2023), who observed an average similarity of 63% ($\sigma$ = 42%) between generated and submitted code, albeit using the Jaccard similarity measure.
@@ 3.3.2. Preceding Activity
To understand when participants decided to use the chatbot, we analyzed the occurrences of prompt types over time in each participant’s session. Figure 7 overviews the prompt type distribution. Additionally, we reviewed the number of submission attempts and the elapsed time before the first occurrence of each prompt type, which are shown in Figure 8.
Figure 7. Timeline showing the distribution of prompts throughout each participant’s experiment session.
Figure 8. Submission attempts (left) and elapsed time (right) preceding the first occurrence of each prompt type.
@@ 3.3.3. Prompting Behavioral Patterns
Figure 9 presents an example of the interaction logs we extracted from participant sessions. We used sequence pattern mining to identify common sequences of actions in these logs. The most common patterns are reported in Figure 10 (left). We report both the total number of occurrences of each pattern, as well as how many participants have followed this pattern at least once during their session. The results show frequent cycles of incorrect submissions, followed by code generation prompts or clipboard events, followed again by incorrect submissions.
We found that while the most common sequences were present in the majority of interaction logs, some had comparable occurrence counts while being exhibited by far fewer participants. This is in part due to most of those patterns containing chatbot-to-quiz clipboard events (i.e. chatbot-to-coderunner), which were less common among participants overall. None of the sequences contained interaction items for a correct submission. We presume this to be because students only needed one correct submission to pass, but made many incorrect submissions prior, thereby overshadowing any possible patterns with correct submissions.
Finally, we additionally performed the same analysis while differentiating submissions by whether they constitute a close match as per section 3.3.1. We report these results in figure 10 (right).
Figure 10. Most common closed interaction subsequences among participants, across all submissions (left) and differentiated by close match status (right).
@@ 4. Discussion
4.1. Summary of Findings
The mixed responses on self-reported usage and the significant fraction of participants who did not engage with the chatbot during the experiment indicate that tools like ChatGPT do not have universal straight recognition or acceptance among students. All participants were made aware of the support chatbot at the beginning of their experiment session. Despite this, usage among those who reported no prior experience with tools was low. We speculate that they may have been hesitant to try it for the first time in this environment, or they may have heard poor experience reports from peers that acted as a deterrent. Interestingly, students who reported regularly using ChatGPT-like tools had more frequent interactions with the chatbot. However, our results clearly indicate that using a chatbot (without prior guidance) for programming task requiring thinking correlates with longer time to solve the task while no impact on task success can be observed.
4.1.1. The Vicious Prompting Cycle: Fix, Apology & Explain, Copy & Refix
The experiment results show students use the chatbot for general knowledge retrieval as well as to get exercises solved (or partially solved) for them. We identified only eight Explain prompts (6.8 %) across all chat histories. Of the Explain prompts we found, four asked “what is wrong with this code?” after having received a compilation error, though without relaying the error in question. Only in two prompts did a participant ask for an explanation of the generated answer after discovering it was incorrect.
A few factors may have contributed to this. Most significantly, unless explicitly instructed otherwise, GPT-3.5 has a tendency to be verbose (Kabir et al., 2024) and confident. We observed that for 97.1% of the generated solutions, an explanation was already provided alongside it. This may have been perceived by students as sufficient, reducing their need to explicitly ask for elaboration. However, aside from compilation errors, none of the participants asked the chatbot to explain why their own solution wasn’t working. In addition, either by previous experience or through their earlier prompts in the study, some participants may have gleaned that the responses from the chatbot were unreliable, and therefore concluded that they had little to gain from its explanations. The comparatively high prevalence of Inform requests, however, indicates some level of trust in the ’s output, at least for general knowledge retrieval tasks.
While reviewing the students’ interactions, we noted that multiple students ended up in cycles of seeing an error message, prompting the chatbot for assistance, implementing the chatbot’s suggestion, and receiving another error message. One such case is presented in Figure 9. Here, the interaction timeline shows an accelerating back-and-forth between prompts and submissions. Through our sequence analysis, we were able to quantify this pattern, which turned out to be the most common interaction pattern across all . As the GPT-3.5 model was unable to produce a correct solution to problem P1 in most cases, even in response to Fix prompts, this was a notably ineffective strategy. Nonetheless, some students appeared intent on ”coercing” the model into generating a correct submission through repeated Codegen prompts. This suggests they were unable to recognize when a tool is incapable of providing effective assistance.
Explanations of potential errors in an automated system have been shown to increase trust and reliance on the system, largely mitigating the negative effect of observing an error (Dzindolet et al., 2003). We observed that ChatGPT often apologized and occasionally even justified its mistakes before attempting to correct itself, which may have increased the students’ trust in the system. This happened in 10 experiment sessions, in fact up to 14 times in one single session (participant 28X).
Additionally, the repeated unmodified submission of generated code indicates these participants did not critically review the model output between generation and submission, even after being made aware of the tool’s limitations. Using for code generation shifts the developer’s focus from ideation (Wei et al., 2024) to parsing and debugging unfamiliar code, which prior work has shown to be a challenge for professional developers (Sarkar et al., 2022; Vaithilingam et al., 2022). However, if students do not understand the generated code they submit, their usage is far more likely to be perceived as academic misconduct (Prather et al., 2023a).
Some participants began their chatbot conversations with questions related to the exercise, but without providing the necessary context. In these cases, the chatbot would then ask for more information from the user. Providing additional context has been shown to significantly improve model output on code generation tasks (Shin et al., 2023). Presumably, participants omitting context were under the impression that the chatbot they were interacting with had already been specifically instructed on their task. Additionally, we observed participants asking for a fix to an incorrect solution without describing the problem, i.e. prompting to the effect of ”that doesn’t work, try again”. More often, however, participants would repeat the problem statement instead, or slightly rephrase it. One approach we observed was to repeat only the instruction itself, omitting the introduction text and context code, potentially in an attempt to “remind” the model of the task at hand. Notably, this behavior has been found to be uncommon among developers at large (Jin et al., 2024).
4.1.2. Straight Usage of GAI Could Lead to a Generation of Lazy Thinking Programmers

### v2

@@ 2.2.2. Experimental Task
For the purposes of generating meaningful student-chatbot interactions, it was important that the task could not be reliably solved by ChatGPT. We found that at the experience level we targeted, isolated code authoring exercises, including most of the existing SE1 course materials, were unsuitable for this reason. Prior work found ChatGPT to struggle with identifying issues in code, especially for logic errors (Tian et al., 2024; Hellas et al., 2023). We thus designed a code reading and comprehension exercise that would be particularly challenging for the , which we confirmed in initial testing. Still, we did not increase the complexity or context size beyond what the students were accustomed to.
The exercise is presented in Figure 2. It consisted of two problems. In P1, participants should identify and trigger a null-related bug in a provided class. Crucially, in our problem setup, the “obvious” solution of passing null wherever possible was insufficient, such that students could not solve the exercise by simply applying a recently learned pattern without tracing the provided code. Instead, they had to combine their knowledge of multiple concepts, such as object state, null dereferencing and operator short-circuiting. Additionally, purposefully producing failure states was a novel task for the students who had so far only learned why they occur and how to prevent them. In P2, participants were then asked to add the missing null check to the original code, to prevent such errors. During the experiment, participants were not allowed to reference other course material, look up information online, or ask other students for help.
The experiment was conducted in-person in a controlled environment between November and December 2023. All participants were students of the same course. To accommodate their schedules, students could participate immediately following each of their respective lab sessions. Students were informed of the study during a course lecture as well as through a Moodle announcement. Additionally, students were approached for recruitment in-person during the lab sessions. As incentive, participants who completed the experiment were able to participate in a gift card raffle.
Participation was completely voluntary and did not have any impact on the actual course grading. We informed participants that they were required to either complete the exercise or spend at least 20 minutes on it. Otherwise, they would not be eligible for the reward and their data would be discarded. There was no upper time limit. While participants were not misled about the purpose of the study, it was kept intentionally vague to reduce potential subject expectancy effects. It was described to them as a study of tool support for learning and problem-solving.
@@ 3. Results
We present the results along the three research questions.
@@ 3.1. Task-Solving Evaluation (RQ1)
Figure 4. Performance in question-answering exercises.
@@ 3.1.1. Coding Exercises
Both GPT models produced syntactically correct code in all responses. Responses from GPT-3.5 had slightly more logical errors than GPT-4, and failed to meet the exercise requirements more than twice as often. Despite not being explicitly instructed to do so, a large majority of the responses were aligned with what a student of the course would already know at that time. Inconsistencies between the described intent and the actual implementation were rare, though in one particularly obvious case, GPT-3.5 provided multiple line comments describing logic that was entirely absent from the code.
Adherence to the course coding conventions was low overall. Neither language model’s baseline code generation – i.e. with no existing code provided as a reference – was in alignment with the formatting guidelines. Even with context code, the majority of generated code did not follow the demonstrated bracketing style, nor were documentation comments added consistently.
@@ 3.2. Chatbot Use (RQ2)
3.2.1. Exercise Completion
In total, 42 students participated in the experiment. Three students participated in an initial test run with an older version of the exercise. Their results were not included in the analysis. Additionally, two students ended the experiment after a short period (4.5 min. and 9 min.) without completing the tasks. We thus concluded that they dropped the participation and discarded their data. This leaves 37 participants for the analysis.
62.2 % of participants completed the entire exercise correctly, and 78.4 % successfully solved at least one problem. The average duration of experiment sessions was 941.9s (M = 859s, $\sigma$ = 559.2s) for students who completed both problems, and 1916.9s (M = 1784s, $\sigma$ = 493.7s) for those who did not. Across the experiment, 556 submission attempts were recorded. The vast majority (89%) were for problem P1. The participants made an average of 13.38 submission attempts for P1 (M = 12, $\sigma$ = 9.5), of which 67.5% resulted in compilation errors and 22.4% in a test failure because their code ran without errors. For P2, participants made 1.65 attempts (M = 1, $\sigma$ = 2.32) on average, of which 27.9% resulted in compilation errors and 13.1% failed because their code did not handle the error case.
Before starting the exercise, we asked participants to rate how challenging they found the course on a 5-point scale, with 1 being very easy and 5 being very difficult. We observed a moderate negative correlation between perceived course difficulty and exercise completion ($r=-0.47,p=0.003$), as well as a moderate positive correlation between perceived difficulty and experiment duration ($r=0.4,p=0.013$). This indicates that the students’ performance in the experiment aligned with their general experience in the course.
3.2.2. Usage of
Before the exercise, we asked participants how often they use tools such as ChatGPT or GitHub Copilot. 16.2% reported to have never used such tools before. 48.6% reported using tools at least monthly (monthly through weekly or more than once a week). We observed a moderate positive correlation between the reported use of tools and the frequency of chatbot interactions during the experiment ($r=0.35,p=0.03$). We did not observe any significant relationship between reported use and performance on the exercise.
@@ 3.2.2. Usage of
The average experiment duration varied significantly between students who did and did not use the chatbot ($t=3.3,p=0.003$). For , the average duration was 26:15 (M = 26:03, $\sigma$ = 10:38). For those who did not use the chatbot, this was 14:36 (M = 10:30, $\sigma$ = 10:27). To account for participants leaving at arbitrary times if they could not solve the exercise after the first 20 minutes, we additionally separated them by exercise completion. When comparing the duration only among completed attempts, the difference is not significant ($t=1.9,p>0.05$). We therefore did not find a relationship between chatbot use and the time to arrive at the correct solution.
@@ 3.2.3. Student Prompts
In this section, all values are relative to the participants who used the chatbot during the experiment. Figure 5 shows the total number of prompts for each type, as well as the number of participants that have submitted such a prompt at least once. Overall, we found that students submitted more Codegen prompts than Support, with a ratio of 62% to 38% [^1] Others prompts are excluded from the totals here, as they were exclusively either off-topic or erroneously submitted., confirming observations by Mailach et al. (2025).
Figure 6 shows the number of occurrences of each prompt type per-participant. We observed Solve prompts from most of the participants, as 73.9% submitted a Solve prompt at some point during the experiment. Interestingly, we also found that some participants submitted more than two Solve prompts. Overall, the most common prompt type was Fix, though this was driven in part by two outliers, who each submitted 13 Fix prompts during their session. The least common prompt type was Explain, observed in only 39.1% of conversations, constituting 6.8% of prompts. Of the 12 Explain prompts we found, five asked “what is wrong with this code?” after encountering an error, though without including the error message in the prompt. Only in two prompts did a participant ask for an explanation of the generated answer after discovering it was incorrect.
Some participants started their chatbot conversations with questions related to the exercise, but without providing the necessary context. In these cases, the chatbot would then ask for more information from the user. Presumably, participants omitting context were under the impression that the chatbot had already been specifically instructed on their task. Additionally, we observed participants prompting for a Fix without describing the problem, i.e. prompting to the effect of “that doesn’t work, try again”. This supports prior findings by Mailach et al. (2025) and Kruse et al. (2024) that some students do not understand what information precisely needs to be shared with the chatbot. More often, however, participants would repeat the problem statement instead, or slightly rephrase it. One approach we observed was to repeat only the instruction itself, omitting the introduction text and context code, potentially in an attempt to “remind” the model of the task at hand. Notably, this behavior seems to be uncommon among developers at large (Jin et al., 2024).
Figure 5. Total number of prompts by type (top) and number of participants that submitted at least one prompt of that type (bottom). Opacity indicates self-reported usage.
@@
Table 4. Correlation analysis of prompt types to preceding survey responses using Kendall’s $\tau$. Statistically significant values are bolded. For individual labels (top rows), significance was adjusted by Bonferroni correction ($\alpha=0.05/5=0.01$).
@@ 3.2.4. Bot Responses
61% of the 172 chatbot responses contained an auto-generated solution to one of the exercise problems, of which 97.1% also contained a written explanation of the solution attempt. In total, 83.1% of responses included an explanation, either for a generated solution, for an example code snippet, or for a general concept.
As intended by the study design, the GPT-3.5 model was unable to solve problem P1 reliably. Out of 80 generated solutions, only 8.8% were correct. The model performed better on P2, where 76.9% of 13 generated solutions were correct.
88.1% of the Inform responses were correct, indicating that when requesting only general information about basic programming concepts, GPT-3.5 does not appear to “hallucinate”. These included questions about when null-related errors occur in Java and how to produce one. Notably, one participant requested the documentation for the class NullPointerException, and the generated response correctly relayed it verbatim.
@@ 3.3.1. Similarity of Consecutive Submissions
Among submissions where participants had at least one Codegen response available to them, the average similarity to the generated code was 0.65 on P1 (M = 0.68, $\sigma$ = 0.26) and 0.66 on P2 (M = 0.71, $\sigma$ = 0.29). These numbers align with findings by Kazemitabaar et al. (2023), who observed an average similarity of 63% ($\sigma$ = 42%) between generated and submitted code, albeit using the Jaccard similarity measure.
@@ 3.3.2. Preceding Activity
To understand when participants decided to use the chatbot, we analyzed the occurrences of prompt types over time in each experiment session. Figure 7 overviews the prompt type distribution. Additionally, we reviewed the number of submission attempts and the elapsed time before the first occurrence of each prompt type, which are shown in Figure 8.
Figure 7. Timeline showing the distribution of prompts throughout each experiment session.
Figure 8. Submission attempts (left) and elapsed time (right) before the first occurrence of a prompt type.
@@ 3.3.3. Prompting Behavioral Patterns
Figure 9 presents an example of the interaction logs we extracted from participant sessions. We used sequence pattern mining to identify common sequences of actions in these logs. The most common patterns are reported in Figure 10 (left). We report both the total number of occurrences of each pattern, as well as how many participants have followed this pattern at least once during their session. The results show frequent cycles of incorrect submissions, followed by code generation prompts or clipboard events, followed again by incorrect submissions. We additionally performed the same analysis while differentiating submissions by whether they constitute a close match as per section 3.3.1. We report these results in Figure 10 (right).
We found that while the most common sequences were present in the majority of interaction logs, some had comparable occurrence counts while being exhibited by far fewer participants. None of the sequences contained interaction items for a correct submission. We assume that any possible patterns with correct submissions were overshadowed, as students only needed one correct submission to pass, but made many incorrect submissions prior.
Figure 10. Most common closed interaction subsequences among participants, with all submissions as the same interaction type (left) and with submissions differentiated by whether they are a close match (right).
@@ 4. Discussion
4.1. Findings and Implications
4.1.1. in Programming Education and Assessment
We found that a student could not pass our course by relying solely on , but almost exclusively due to secondary assessments that go beyond code correctness and which require demonstration of conceptual understanding in a controlled environment. Both GPT models were able to generate correct code and essay answers for most of the exercises. In a non-personal setting, such as remote homework submissions, a student could therefore likely pass the course using GPT-4, even with manually graded assignments. Though the occasionally missed individual requirements, such mistakes could also arise through honest misunderstandings by students. They would be easily corrected based on tutor feedback. In almost all cases, the GPT output is thus nearly indistinguishable from that of a well-performing student. This renders submission-based assignments significantly less effective for assessing students’ understanding of submitted code or underlying concepts, despite the latter being a key requirement among educators to consider use acceptable (Prather et al., 2023a).
However, with the format of our course, minor mistakes can become a significant challenge for a student presenting AI-generated work as their own. Because the necessary corrections are also minor, the students—having presumably put significant thought into their answer already—are expected to apply them on the spot. Inability to do so, or to at least discuss potential approaches with the tutor, would be a strong indicator of academic dishonesty. Additionally, our students must be able to justify and explain the use of concepts, syntax, or APIs that have not been covered in the course yet. Without a solid understanding of the existing course materials, it would be difficult to even identify the elements in the code that may lead to further questioning, and then convincingly explain them. As this risk is unrelated to the reasoning capabilities of the , it cannot be mitigated by using more advanced models. This observation is supported by the lack of improvement in the Learn category between GPT-3.5 and GPT-4.
The students’ self-reported usage and the engagement with the chatbot in the experiment were mixed. While this indicates that tools like ChatGPT do not yet have universal acceptance among students, we found that self-reported regular usage was associated with more frequent chatbot interactions and a higher propensity for code generation requests. This suggests that as students self-learn and gain experience with , they become more reliant on its code output.
We derive two major recommendations. First, the findings suggest GenAI should be introduced as a tool early in software development courses to avoid misperception through self-learning. Particularly, instructors should highlight various usage strategies beyond generating code, e.g. to reason about and understand code—in combination with other tools like debuggers and visualizations. Limitations and tendencies of current models should be discussed in classes to create awareness, e.g. about hallucinations or overconfidence. Second, we recommend, if possible, to shift towards interview-like assessments and code reviews instead of or in addition to homework. This creates more opportunities to detect academic dishonesty and knowledge gaps without requiring a categorical ban on tools. Future work could explore whether the feedback and the evaluation in these assessments could itself be assisted by , possibly using a tutor-in-the-loop approach to lessen the resource burden on educators.
4.1.2. Learning to Think with and to Develop, Compare, and Challenge Solution Paths
In our experiment, the participants used the chatbot for general knowledge retrieval as well as to have the task solved (or partially solved) for them. Requests for explanations were exceedingly rare, and we did not observe students asking any comprehension questions about the exercise itself, or why their own solution wasn’t working. In contrast, we observed that 97.1% of the generated solutions already included an explanation. This may have been perceived by students as sufficient, reducing their need to explicitly ask for elaboration. Unless explicitly instructed otherwise, GPT-3.5 has a tendency to be verbose (Kabir et al., 2024) and confident (Hellas et al., 2023).